OpenStudy (kingsalo7):

How many atoms are found in 1.56 moles of Na? 9.39 x 1023 atoms 3.76 x 1023 atoms

OpenStudy (sweetburger):

\[1.56moles(Na)\times \frac{ 6.022\times10^\left( 23 \right)atoms(Na) }{ 1 mole(Na) }\]

OpenStudy (sweetburger):

moles of (Na) cancel and you are left with units in terms of atoms. Just complete the multiplication.

OpenStudy (roku):

am lost